# Search for Light Dark Matter with 259 Days of Data in PandaX-4T Minzhen Zhang

### v3

@@ Abstract
We establish the most stringent constraints on spin-independent dark matter-nucleon interactions within a mass range of 2.5-5.0 GeV/, spin-dependent neutron-only interactions within 2.0-4.6 GeV/, and spin-dependent proton-only interactions within 2.0-3.6 GeV/.
@@ 2 GeV/
The results establish the most stringent limits for SI, SD (neutron-only), and SD (proton-only) DM-nucleon scatterings in the mass range of approximately [2.5, 5.0], [2.0, 4.6], and [2.0, 3.6] GeV/c2, respectively.
@@
Our analysis establishes new exclusion limits at 90% CL for: spin-independent DM-nucleon scattering in the mass range [2.5, 5.0] GeV/, spin-dependent DM-nucleon scattering ([2.0, 4.6] GeV/ for neutron-only coupling, and [2.0, 3.6] GeV/ for proton-only coupling), and LDM-electron scattering (from 100 MeV/ to 10 GeV/).

### v4

@@ Abstract
We establish the most stringent constraints on spin-independent dark matter-nucleon interactions within a mass range of 2.5-5.0 GeV/, spin-dependent neutron-only interactions within 2.0-5.3 GeV/, and spin-dependent proton-only interactions within 2.0-3.8 GeV/.
@@ 2 GeV/
The results establish the most stringent limits for SI, SD (neutron-only), and SD (proton-only) DM-nucleon scatterings in the mass range of approximately [2.5, 5.0], [2.0, 5.3], and [2.0, 3.8] GeV/c2, respectively.
@@
Our analysis establishes new exclusion limits at 90% CL for: spin-independent DM-nucleon scattering in the mass range [2.5, 5.0] GeV/, spin-dependent DM-nucleon scattering ([2.0, 5.3] GeV/ for neutron-only coupling, and [2.0, 3.8] GeV/ for proton-only coupling), and LDM-electron scattering (from 100 MeV/ to 10 GeV/).